# Two-dimensional inference of divertor plasma characteristics: advancements to a multi-instrument Bayesian analysis system

### v1

@@ A.2 Helium Emission Models
MAST-U’s MWI and TCV’s MANTIS cameras also routinely capture Helium-I emission lines which has been shown [40] to provide information on $n_{\mathrm{e}}$ and $T_{\mathrm{e}}$. The nature of integrated data analysis is such that a single inaccurate forward model can offset the overall inference. As such we explicitly include parameters for $n_{{\mathrm{He}}^{{0}}}$ and $n_{{\mathrm{He}}^{{1+}}}$ at each mesh vertex (at a cost of 2V additional free parameters). Doing so avoids assumptions of ionisation balance models and permits inclusion of both electron-impact excitation (of $\mathrm{He}^{0}$) and electron-ion recombination (of $\mathrm{He}^{+}$) contributions to emission. He I PECs were provided by ADAS [22] and implemented as with (5).
In conditions with significant hydrogen atomic densities, charge-exchange between helium and hydrogen could occur. This can change both the ionisation fractions of $\mathrm{He}$ (which is included in our model through the $n_{{\mathrm{He}}^{{0}}}$ and $n_{{\mathrm{He}}^{{1+}}}$ parameters) and lead to the emission of He I [57]. Although this is neglected in our emission model, since population coefficients for this interaction are not freely available, having neutral hydrogen density, $n_{\mathrm{H}}$, as a parameter in our model means that charge exchange can trivially be included once population coefficients for this interaction become available.
The 668, 728 and 502 nm singlet He I lines were chosen for the helium forward models. This followed from the work by [43], which used these transitions due to their insensitivity to transport of metastable states, which can impact singlet to triplet He I line ratios. Detailed studies on TCV using He I emission [40] have suggested that the magnetic field as well as molecular effects could impact He I emission. However, based on those results, and the relatively low magnetic field at MAST-U (0.7 T on-axis), we would not expect that this would impact the chosen He I singlet transitions significantly.
@@ B.1 Challenges associated with Q-mol inferences and possible improvements
We have observed that there is a negative correlation in the samples of $n_{\mathrm{H}}$ and $Q_{\mathrm{mol}}$ (Figure 3 c, d). As shown in (7), both $n_{\mathrm{H}}$ and $Q_{\mathrm{mol}}$ can be freely altered to match the required (i.e., synthetically measured) emissivity. This suggests that the emission problem is ill-posed for these parameters, which is exacerbated by the lack of strong priors for $Q_{\mathrm{mol}}$.
@@
The inferences of $Q_{\mathrm{mol}}$ could be improved through: introducing priors on the $Q_{\mathrm{mol}}$ profile; or the inclusion of additional diagnostics that provide information on $n_{\mathrm{H}}$ and/or $Q_{\mathrm{mol}}$. This information can be provided indirectly through techniques that enable separating EIE and PMI hydrogen emission. One possibility is to utilise the $D_{2}$ Fulcher emission that has been successfully applied to line-integrated spectroscopy [27]. An additional Balmer-line camera in the upstream region, such as the X-Point imaging camera recently installed on MAST-U, could be incorporated into the IDA to improve the inferences in the region where the MWI coverage is limited.
@@ B.2 Numerical complexities of characterising the Posterior Distribution
The nature of our posterior distribution posed three key challenges: the number of parameters, the correlations between parameters and multi-modality. These challenges especially hamper adequate sampling of the posterior distribution. Hamiltonian Monte Carlo (HMC) [37], with a mass matrix constructed from an approximation of the Hessian in the MAP vicinity countered these challenges. However, significant computational resources were required due to the small time-step necessary to avoid divergence in the HMC algorithm and the accompanying large number of leap-frog steps necessary to avoid inefficient random-walk like behaviour. This is observed in this work in figure 3 where only 63% of $n_{\mathrm{H}}$ and $n_{{\mathrm{He}}^{{1+}}}$ parameters contain the true value within their 95% HDI thus demonstrating incomplete sampling.
The finding of the MAP, using the genetic algorithm outlined in the appendix of [23], was negatively impacted by the use of the Cauchy prior probability function. The heavier tails, in comparison to the Gaussian distribution, made for less-pronounced peaks in the probability distribution. This increased the number of local-maxima far from the global maximum resulting in a longer time for the genetic algorithm to reach convergence. This was somewhat countered by the hybrid use of the momentum-based, first-order adam optimiser [58] and the second-order L-BFGS optimiser [59] which carry different benefits in different regions of the parameter space.
@@ B.3 Improving the analysis through spatially-dependent priors and its limitations
Figure 6 of [23] outlined how varying the ‘strength’ (the reciprocal of the $\sigma$ parameter for a Gaussian distribution) of the smoothing prior probability distribution can significantly alter the inference. Investigations found that each plasma state (i.e. the detached case and the attached case) had different optimal smoothing strengths that cannot be known a priori.
@@
The choice of a Cauchy distribution to handle the smoothing priors came with limitations. Since it did not penalise non-smooth parameter combinations as harshly as a Gaussian distribution, the optimiser was slower to converge on a smooth solution. Additionally, diagnostic noise could suggest a sharp feature in the plasma state which a Cauchy smoothing prior could incorrectly permit. This can be seen by the un-smooth drop in electron temperature in the region of (1.2 -1.7)m of figures 2 and 8. Consequently, when moving to experimental data, a more sophisticated approach to handling the crucial smoothing priors may be required.
@@ C.2 Prior details
C.3 Spatially-Independent Priors
@@
Prior ($j$)
Argument ($f_{j}(\theta_{k})$)
$\gamma_{j}$
Justification
@@
$\max\left(\frac{P^{k}_{e}}{P_{e}^{\mathrm{max}}}-1,0\right)$
${1.5}\times 10^{-2}$
Static electron pressure is not expected to exceed $P_{e}^{max}=7\times 10^{20}eVm^{-3}$ in MAST-U Super-X divertor, limit found through analysis of multiple SOLPS-ITER cases (figure 4 of [23]).
@@
$\quad f_{n_{0}}^{k}=\frac{n_{0}^{k}}{n_{e}^{k}+n_{0}^{k}}$
$\max\left(\frac{f_{0}}{f_{n_{0}}^{k}}-1,\frac{f_{n_{0}}^{k}}{f(T_{e})}-1,0\right)$
(${2.3}\times 10^{-3}$,
@@
In MAST-U Super-X divertor, neutral fraction is not expected to exceed, f(T_e) = (1-c)exp(-Tekl)+c, where $c=0.04$ and $l=5eV$. Similarly, the fraction was expected to remain above $f_{0}=0.002$. Limits found through analysis of multiple SOLPS-ITER cases (figure 4 of [23]).
@@
$\max\left(\frac{I_{0}(T_{e})}{I_{He^{1+}}^{k}}-1,\frac{I_{He^{1+}}^{k}}{I_{1}(T_{e})}-1,0\right)$
(${1.4}\times 10^{-3}$,
@@
$\max\left(\frac{C_{0}}{C_{He}^{k}}-1,\frac{C_{He}^{k}}{C_{1}}-1,0\right)$
(${3.6}\times 10^{-2}$,
@@
$\max\left(\frac{\log(A)}{\log(\theta_{F}^{k})}-1,\frac{\log(\theta_{F}^{k})}{\log(B)}-1,0\right)$
$A\in($0.2, ${1}\times 10^{16}$, ${1}\times 10^{15}$, ${1}\times 10^{-3}$, ${1}\times 10^{15}$, ${1}\times 10^{13}$)))) respectively.
@@ C.4 Spatially-Dependent Priors
Prior ($j$)
Argument ($f_{j}(\theta_{k},...)$)
$\gamma_{j}$
Justification
@@
From the two point model, expected minimally factor two reduction in static electron pressure from upstream to target along a flux tube[60]. Since this pressure reduction argument loses validity if the upstream mesh vertex (‘$k\mathrm{,upstream}$’ is below the detachment threshold, this prior is only applied to flux tubes with an upstream electron temperature above 5eV. Prior is applied only at mesh vertices along the target.
@@
$\max\left(0,\partial_{s_{\parallel}}F^{k}\right)$
(${2.4}\times 10^{-3}$,
@@
$\max\left(0,\partial_{s_{\parallel}}Q^{k}_{\parallel}\right)$
${4.0}\times 10^{+2}$
Expected monotonic reduction in parallel heat flux on a path to the target, $s_{\parallel}$, parallel to surfaces of constant magnetic flux. The prior’s argument form means the thermal conductivity, $\kappa$, can take any positive value.
@@
$\partial^{2}_{s_{\parallel}s_{\parallel}}\log\left(F^{k}\right)$
(${1.2}\times 10^{-1}$,
@@
    $F^{k}\in(n_{0}^{k},n_{He^{0}}^{k})$
$\partial^{2}_{s_{\theta,\parallel}s_{\theta,\parallel}}\log\left(F^{k}\right)$
(${1.9}\times 10^{+1}$,
@@
$(T_{e}^{k},n_{e}^{k},n_{0}^{k},n_{He^{0}}^{k},n_{He^{1+}}^{k})$
$\partial^{2}_{s_{\psi_{N}}s_{\psi_{N}}}\log\left(F^{k}\right)$
(${1.8}\times 10^{+2}$,

### v2

@@ A.2 Helium Emission Models
MAST-U’s MWI and TCV’s MANTIS cameras also routinely capture Helium-I emission lines which has been shown [Linehan2023] to provide information on $n_{\mathrm{e}}$ and $T_{\mathrm{e}}$. The nature of integrated data analysis is such that a single inaccurate forward model can offset the overall inference. As such we explicitly include parameters for $n_{{\mathrm{He}}^{{0}}}$ and $n_{{\mathrm{He}}^{{1+}}}$ at each mesh vertex (at a cost of 2V additional free parameters). Doing so avoids assumptions of ionisation balance models and permits inclusion of both electron-impact excitation (of $\mathrm{He}^{0}$) and electron-ion recombination (of $\mathrm{He}^{+}$) contributions to emission. He I PECs were provided by ADAS [adas] and implemented as with equation (5).
In conditions with significant hydrogen atomic densities, charge-exchange between helium and hydrogen could occur. This can change both the ionisation fractions of $\mathrm{He}$ (which is included in our model through the $n_{{\mathrm{He}}^{{0}}}$ and $n_{{\mathrm{He}}^{{1+}}}$ parameters) and lead to the emission of He I [Munoz2019]. Although this is neglected in our emission model, since population coefficients for this interaction are not freely available, having neutral hydrogen density, $n_{\mathrm{H}}$, as a parameter in our model means that charge exchange can trivially be included once population coefficients for this interaction become available.
The 668, 728 and 502 nm singlet He I lines were chosen for the helium forward models. This followed from the work by [Biggelaar2022], which used these transitions due to their insensitivity to transport of metastable states, which can impact singlet to triplet He I line ratios. Detailed studies on TCV using He I emission [Linehan2023] have suggested that the magnetic field as well as molecular effects could impact He I emission. However, based on those results, and the relatively low magnetic field at MAST-U (0.7 T on-axis), we would not expect that this would impact the chosen He I singlet transitions significantly.
@@ B.1 Challenges associated with Q-mol inferences and possible improvements
We have observed that there is a negative correlation in the samples of $n_{\mathrm{H}}$ and $Q_{\mathrm{mol}}$ (Figure 3 c, d). As shown in equation (7), both $n_{\mathrm{H}}$ and $Q_{\mathrm{mol}}$ can be freely altered to match the required (i.e., synthetically measured) emissivity. This suggests that the emission problem is ill-posed for these parameters, which is exacerbated by the lack of strong priors for $Q_{\mathrm{mol}}$.
@@
The inferences of $Q_{\mathrm{mol}}$ could be improved through: introducing priors on the $Q_{\mathrm{mol}}$ profile; or the inclusion of additional diagnostics that provide information on $n_{\mathrm{H}}$ and/or $Q_{\mathrm{mol}}$. This information can be provided indirectly through techniques that enable separating EIE and PMI hydrogen emission. One possibility is to utilise the $D_{2}$ Fulcher emission that has been successfully applied to line-integrated spectroscopy [Verhaegh2023c]. An additional Balmer-line camera in the upstream region, such as the X-Point imaging camera recently installed on MAST-U, could be incorporated into the IDA to improve the inferences in the region where the MWI coverage is limited.
@@ B.2 Numerical complexities of characterising the Posterior Distribution
The nature of our posterior distribution posed three key challenges: the number of parameters, the correlations between parameters and multi-modality. These challenges especially hamper adequate sampling of the posterior distribution. Hamiltonian Monte Carlo (HMC) [neal2011mcmc], with a mass matrix constructed from an approximation of the Hessian in the MAP vicinity countered these challenges. Figure 9 displays the error in our inferred parameters (a) and the inferred uncertainty (b). We find the regions of larger error do indeed correspond to increased inferred uncertainty. This suggests that, despite the complex posterior distribution, HMC was able to adequately characterise it.
The HMC algorithm required significant computational resources. This was due to the small time-step necessary to avoid divergence in the HMC algorithm and the accompanying large number of leap-frog steps necessary to avoid inefficient random-walk like behaviour. This is observed in this work in figure 3 where only 63% of $n_{\mathrm{H}}$ and $n_{{\mathrm{He}}^{{1+}}}$ parameters contain the true value within their 95% HDI suggesting incomplete sampling. An improved method for finding the mass matrix used in the HMC algorithm is required for more efficient sampling.
The finding of the MAP, using the genetic algorithm outlined in the appendix of [Bowman_2020], was negatively impacted by the use of the Cauchy prior probability function. The heavier tails, in comparison to the Gaussian distribution, made for less-pronounced peaks in the probability distribution. This increased the number of local-maxima far from the global maximum resulting in a longer time for the genetic algorithm to reach convergence. This was somewhat countered by the hybrid use of the momentum-based, first-order adam optimiser [kingma2017adam] and the second-order L-BFGS optimiser [Liu1989] which carry different benefits in different regions of the parameter space.
@@ B.3 Improving the analysis through spatially-dependent priors and its limitations
Figure 6 of [Bowman_2020] outlined how varying the ‘strength’ (the reciprocal of the $\sigma$ parameter for a Gaussian distribution) of the smoothing prior probability distribution can significantly alter the inference. Investigations found that each plasma state (i.e. the detached case and the attached case) had different optimal smoothing strengths that cannot be known a priori.
@@
The choice of a Cauchy distribution to handle the smoothing priors came with limitations. Diagnostic noise could suggest a sharp feature in the plasma state which a Cauchy smoothing prior could incorrectly permit. This can be seen by the un-smooth drop in electron temperature in the region of (1.2 -1.7)m of figures 2 and 7. Consequently, when moving to experimental data, a more sophisticated approach to handling the crucial smoothing priors may be required.
Figure 9: Visualisation of error (a) and uncertainty (b) for the detached case using posterior (i).
@@ C.2 Prior details
Details of the spatially-independent priors are given in Table 1 and details of the spatially-dependent priors are given in Table 2.
@@
Prior ($j$) Argument ($f_{j}(\theta_{k})$) $\gamma_{j}$ Justification
@@
$\max\left(\frac{P^{k}_{e}}{P_{e}^{\mathrm{max}}}-1,0\right)$ ${1.5}\times 10^{-2}$ Static electron pressure is not expected to exceed $P_{e}^{max}=7\times 10^{20}eVm^{-3}$ in MAST-U Super-X divertor, limit found through analysis of multiple SOLPS-ITER cases (Figure 4 of [Bowman_2020]).
@@
$\quad f_{n_{H}}^{k}=\frac{n_{H}^{k}}{n_{e}^{k}+n_{H}^{k}}$
$\max\left(\frac{f_{0}}{f_{n_{H}}^{k}}-1,\frac{f_{n_{H}}^{k}}{f(T_{e})}-1,0\right)$ (${2.3}\times 10^{-3}$,
@@
In MAST-U Super-X divertor, neutral fraction is not expected to exceed, f(T_e) = (1-c)exp(-T_e^k/l)+c, where $c=0.04$ and $l=5eV$. Similarly, the fraction was expected to remain above $f_{0}=0.002$. Limits found through analysis of multiple SOLPS-ITER cases (Figure 4 of [Bowman_2020]).
@@
$\max\left(\frac{I_{0}(T_{e})}{I_{He^{1+}}^{k}}-1,\frac{I_{He^{1+}}^{k}}{I_{1}(T_{e})}-1,0\right)$ (${1.4}\times 10^{-3}$,
@@
$\max\left(\frac{C_{0}}{C_{He}^{k}}-1,\frac{C_{He}^{k}}{C_{1}}-1,0\right)$ (${3.6}\times 10^{-2}$,
@@
$\max\left(\frac{\log(A)}{\log(\theta_{F}^{k})}-1,\frac{\log(\theta_{F}^{k})}{\log(B)}-1,0\right)$ $A\in($0.2, ${1}\times 10^{16}$, ${1}\times 10^{15}$, ${1}\times 10^{-3}$, ${1}\times 10^{15}$, ${1}\times 10^{13}$)))) respectively.
@@
Prior ($j$) Argument ($f_{j}(\theta_{k},...)$) $\gamma_{j}$ Justification
@@
From the two point model, expected minimally factor two reduction in static electron pressure from upstream to target along a flux tube[stangeby2000plasma]. Since this pressure reduction argument loses validity if the upstream mesh vertex (‘$k\mathrm{,upstream}$’ is below the detachment threshold, this prior is only applied to flux tubes with an upstream electron temperature above 5eV. Prior is applied only at mesh vertices along the target.
@@
$\max\left(0,\partial_{s_{\parallel}}F^{k}\right)$ (${2.4}\times 10^{-3}$,
@@
$\max\left(0,\partial_{s_{\parallel}}Q^{k}_{\parallel}\right)$ ${4.0}\times 10^{+2}$ Expected monotonic reduction in parallel heat flux on a path to the target, $s_{\parallel}$, parallel to surfaces of constant magnetic flux. The prior’s argument form means the thermal conductivity, $\kappa$, can take any positive value.
@@
$\partial^{2}_{s_{\parallel}s_{\parallel}}\log\left(F^{k}\right)$ (${1.2}\times 10^{-1}$,
@@
    $F^{k}\in(n_{H}^{k},n_{He^{0}}^{k})$
$\partial^{2}_{s_{\theta,\parallel}s_{\theta,\parallel}}\log\left(F^{k}\right)$ (${1.9}\times 10^{+1}$,
@@
$(T_{e}^{k},n_{e}^{k},n_{H}^{k},n_{He^{0}}^{k},n_{He^{1+}}^{k})$
$\partial^{2}_{s_{\psi_{N}}s_{\psi_{N}}}\log\left(F^{k}\right)$ (${1.8}\times 10^{+2}$,